Determine the future value of a $5,000 compound-interest Canada Savings Bond at 8.5%/a, compounded annually after each amount of time.

4 years


1
Expert's answer
2021-07-27T09:20:12-0400

Given amount =$5,000

Compound Interest Rate (R) =8.5%

Time (n) =4years

So Future Value ;


"A=P(1 +\frac {R} {100}) ^n"


"A=5,000(1 +\frac {8.5} {100}) ^4"


"A=5000+(1.385)^4"


"A=\$6,929.29"


Therefore the required Future Value is $6,929.29